# 3D Gaussian Ray Tracing: Fast Tracing of Particle Scenes

### v2

@@ Abstract.
Particle-based representations of radiance fields such as 3D Gaussian Splatting, have found great success for reconstructing and re-rendering of complex scenes. Most existing methods render particles via rasterization, projecting them to screen space tiles for processing in a sorted order. This work instead considers ray tracing the particles, building a bounding volume hierarchy and casting a ray for each pixel using high-performance GPU ray tracing hardware. To efficiently handle large numbers of semi-transparent particles, we describe a specialized rendering algorithm which encapsulates particles with bounding meshes to leverage fast ray-triangle intersections, and shades batches of intersections in depth-order. The benefits of ray tracing are well-known in computer graphics: processing incoherent rays for secondary lighting effects such as shadows and reflections, rendering from highly-distorted cameras common in robotics, stochastically sampling rays, and more. With our renderer, this flexibility comes at little cost compared to rasterization. Experiments demonstrate the speed and accuracy of our approach, as well as several applications in computer graphics and vision. We further propose related improvements to basic Gaussian representation, including a simple use of generalized kernel functions which significantly reduces particle hit counts.
@@
submissionid: 536ccs: Computing methodologies Renderingccs: Computing methodologies Reconstruction
Figure 1. We propose a method for fast forward and inverse ray tracing of particle-based scene representations such as Gaussians. The main idea is to construct encapsulating primitives around each particle, and insert them into a BVH to be rendered by a ray tracer specially adapted to the high density of overlapping particles. Efficient ray tracing opens the door to many advanced techniques in computer graphics and vision, including secondary ray effects like mirrors, refractions and shadows, as well as highly-distorted cameras with rolling shutter effects and even stochastic sampling of rays.
Project page: GaussianTracer.github.io
@@ 1. Introduction
Figure 2. Ray Tracing Effects: The ray based nature of our pipeline makes it easily compatible with conventional ray based visual effects at test-time, including reflections (top left), depth of field (top middle), refractions (bottom left), hard-shadows cast by meshes (bottom middle) and combinations thereof (right).
@@ 2.1. Novel-View Synthesis and Neural Radiance Fields
Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020) have revolutionized the field of novel-view synthesis by representing the scenes in terms of a volumetric radiance field encoded in a coordinate-based neural network. This network can be queried at any location to return the volumetric density and view-dependent color. The photo-realistic quality of NeRFs has made them the standard representation for novel-view synthesis. Follow-up works have focused on improving the speed (Müller et al., 2022; Reiser et al., 2021), quality (Barron et al., 2021, 2022, 2023), and surface representation (Wang et al., 2021; Yariv et al., 2021; Li et al., 2023; Wang et al., 2023a). NeRF has also been extended to large-scale scenes (Turki et al., 2022; Li et al., 2024a), sparse inputs views (Niemeyer et al., 2022) and in-the-wild image collections (Martin-Brualla et al., 2021). Finally, several works investigated ways to speed up the inference by baking the neural fields to more performant representations (Reiser et al., 2023; Wang et al., 2023b; Duckworth et al., 2023; Reiser et al., 2024). While achieving high quality and fast rendering speeds, these methods often employ compute-expensive multi-stage training procedures.
@@ 2.2. Point-Based and Particle Rasterization
Jointly, these works have made significant progress, but they still inherit limitations of rasterization. Indeed, they are not able to represent highly distorted cameras, model secondary lighting effects, or sensor properties such as rolling shutter or motion blur. Several works have tried to work around these limitations.  Niemeyer et al. (2024) first train a Zip-NeRF (Barron et al., 2023) that can model distorted and rolling shutter cameras and then render perfect pinhole cameras from the neural field and distill them into a 3DGS representation. This allows for fast inference, but is still limited to perfect pinhole cameras. To model secondary lighting effects Gao et al. (2023) combines rasterization with particle-based ray tracing for the visibility pass. Finally,  Seiskari et al. (2024) model the motion blur and rolling shutter of the camera by approximating them in screen space through rasterization with pixel velocities. Unlike these works, we formulate a principled method for efficient ray tracing of volumetric particles, which natively alleviates all the limitations mentioned above and further allows us to simulate effects such as depth of field and perfect mirrors.
@@ 2.3. Differentiable Ray Tracing of Volumetric Particles
Our formulation takes root in these precursory works. However as opposed to (Knoll et al., 2019), it is guaranteed to process every particle intersecting the ray. And contrary to (Brüll and Grosch, 2020) the hit processing order is consistent, which ensures the differentiability of our tracing algorithm.
Compared to rasterization, differentiable ray tracing of semi-transparent particles has seen much less progress in recent years. Perhaps the most similar rendering formulation to ours was proposed in Fuzzy Metaballs (Keselman and Hebert, 2022, 2023), but it is limited to scenes with a small set of 3D Gaussian particles (several tens) and images with very low resolution. Different to Fuzzy Metaballs, our method can easily handle scenes with several millions of particles from which it can render full HD images in real-time.
@@ 4.1. Bounding Primitives
Figure 4. Proxy Geometries: Examples of BVH primitives considered.
@@
(7) $\bm{v}\leftarrow\bm{v}\sqrt{2\log(\sigma/\alpha_{\textrm{min}})}\bm{S}\bm{R}^{T}+\bm{\mu}.$
@@ 4.2. Ray Tracing Renderer
Our renderer builds on past approaches for tracing semi-transparent surfaces or particles:  Knoll et al. (2019) repeated gather slabs of particles and sort within each slab, while  Brüll and Grosch (2020) process all semi-transparent surfaces into a $k$-buffer, merging adjacent particles when the list overflows. As discussed in 2.3, because of their approximations, these algorithms do not produce a consistent rendering, which prevents differentiation and generates artifacts.
Figure 5. Rendering: on each round of tracing, the next $k$ closest hit particles are collected and sorted into depth order along the ray, the radiance is computed in-order, and the ray is cast again to process the next chunk.
@@
Figure 5, Figure 3, Procedure 4.3, and Procedure 4.3 summarize our approach. To compute incoming radiance along each ray, a ray-gen program traces a ray against the BVH to gather the next $k$ particles, using an any-hit program to maintain a sorted buffer of their indices. For efficiency, at this stage the particle response is not yet evaluated; all primitive hits are treated as intersected particles. The ray-gen program then iterates through the sorted array of primitive hits, retrieves the corresponding particle for each, and renders them according to Equation 6. The process then repeats, tracing a new ray from the last rendered particle to gather the next $k$ particles. The process terminates once all particles intersecting the ray are processed, or early-terminates as soon as enough particle density is intersected to reach a predefined minimum transmittance $T_{\textrm{min}}$. Compared to past approaches this renderer allows for processing the intersection in a consistent order, without missing any particle nor approximating the transmittance.
@@ 4.3. Evaluating Particle Response
5:     if $\mathcal{H}[i].\tau_{\textrm{hit}}<h.\tau_{\textrm{hit}}$ then
@@
4.4. Differentiable Ray Tracing and Optimization
Differentiable Rendering
Beyond forward-rendering of particle scenes, our ray tracing renderer is also differentiable, to support optimizing particle scenes from observed data. To backpropagate (i.e., reverse-mode differentiate) through the renderer with respect to particle parameters, we first perform an ordinary forward-pass render and compute the desired objective functions. Then, in the backward pass we re-cast the same rays to sample the same set of particles in order, computing gradients with respect to each shading expression and accumulating gradients in shared buffers with atomic scatter-add operations. We implemented all derivative expressions by hand in an NVIDIA OptiX ray-gen program which is structurally similar to  Procedure 4.3.
@@ 4.5. Particle Kernel Functions
  • Cosine wave modulations:
    (10) $\hat{\rho}_{c}(\bm{x})=\hat{\rho}(\bm{x})(0.5+0.5cos(\psi\bm{R}^{T}\bm{S}^{-1}(\bm{x}-\bm{\mu})))$
@@ 5. Experiments and Ablations
In the experiments that follow, we will refer to two variants of our method. The Ours (reference) variant corresponds to (Kerbl et al., 2023) as closely as possible. It employs regular 3D Gaussian particles, and leaves the optimization hyperparameters unchanged. We treat this as a high-quality configuration. The Ours variant is adapted based on the experiments that follow, improving runtime speed at a slight expense of quality. It uses degree-2 generalized Gaussian particles, a density learning rate to $0.09$ during optimizing, as well as optimizing with incoherent rays in a batch size of $2^{19}$ starting after 15,000 training iterations.
@@ 5.1. Novel View Synthesis Benchmarks
MipNeRF360 Tanks & Temples Deep Blending
[…]
Plenoxels 23.63 0.670 0.443 2.1GB 21.08 0.719 0.379 2.3GB 23.06 0.795 0.510 2.7GB
INGP-Base 26.43 0.725 0.339 13MB 21.72 0.723 0.330 13MB 23.62 0.797 0.423 13MB
INGP-Big 26.75 0.751 0.299 48MB 21.92 0.745 0.305 48MB 24.96 0.817 0.390 48MB
MipNeRF360 29.23 0.844 0.207 8.6MB 22.22 0.759 0.257 8.6MB 29.40 0.901 0.245 8.6MB
3DGS (paper) 28.69 0.870 0.182 734MB 23.14 0.841 0.183 411MB 29.41 0.903 0.243 676MB
3DGS (checkpoint) 28.83 0.867 0.228 763MB 23.35 0.837 0.222 422MB 29.43 0.898 0.314 704MB
Ours (reference) 28.88 0.871 0.179 387MB 23.03 0.853 0.158 519MB 29.89 0.908 0.234 329MB
Ours 28.71 0.854 0.202 383MB 23.20 0.830 0.186 489MB 29.23 0.900 0.242 287MB
@@ 5.1.1. MipNeRF360
Table 1 shows quantitative results, while novel-views are qualitatively compared in Figure 7. In terms of quality, our method performs on par or slightly better than 3DGS (Kerbl et al., 2023) and other state-of-the-art methods.
For this dataset, we also compare our method against the recent top-tier method of Zip-NeRF (Barron et al., 2023) which achieves 30.38 PSNR. In terms of runtime (Table 2), at 78 FPS our efficient ray tracing implementation is approximately three times slower than rasterization (238 FPS), while maintaining interactive speeds compared to high-quality ray-marching based works MipNeRF360 and Zip-NeRF (¡1 FPS).
@@ 5.2.2. Tracing Algorithms
  • Slab tracing (Knoll et al., 2019) (SLAB): tracing one ray per slab, order independent collection of the $k$-first hits in the any-hit program, sorting and integrating the hits in the ray-gen program.
@@
For each algorithms with a choice of parameters (size of the array, number of slabs, or number of samples), we perform a parameter sweep and present the best-performing setting.
@@ 6.1. Ray-Based Effects
First, we the radiance field rendering pipeline with a variety of visual effects which are naturally compatible with ray tracing (Figure 2 and  12). Here we demonstrate only manually-specified forward rendering, although inverse rendering in concert with these effects is indeed supported by our approach, and is a promising area for ongoing work.
@@
Optical ray effects are supported by interleaved tracing of triangular faces and Gaussian particles. Precisely, we maintain an extra acceleration structure consisting only of mesh faces for additional inserted geometry. When casting each ray in Procedure 4.3, we first cast rays against inserted meshes; if a mesh is hit, we render all particles only up to the hit, and then compute a response based on the material. For refractions and reflections, this means continuing tracing along a new redirected ray according to the laws of optics. For non-transparent diffused meshes, we compute the color and blend it with the current ray radiance, then terminate the ray.
@@
Following (Müller et al., 2022), we simulate depth of field by progressively tracing multiple independent ray samples per pixel (spp), weighted together with a moving average to denoise the output image. The examples in Figures  2 and 12 use 64-256 spp, although convergence is often reached with fewer samples by selecting subsamples with low discrepancy sequences (Burley, 2020).

### v3

@@ Abstract.
Particle-based representations of radiance fields such as 3D Gaussian Splatting have found great success for reconstructing and re-rendering of complex scenes. Most existing methods render particles via rasterization, projecting them to screen space tiles for processing in a sorted order. This work instead considers ray tracing the particles, building a bounding volume hierarchy and casting a ray for each pixel using high-performance GPU ray tracing hardware. To efficiently handle large numbers of semi-transparent particles, we describe a specialized rendering algorithm which encapsulates particles with bounding meshes to leverage fast ray-triangle intersections, and shades batches of intersections in depth-order. The benefits of ray tracing are well-known in computer graphics: processing incoherent rays for secondary lighting effects such as shadows and reflections, rendering from highly-distorted cameras common in robotics, stochastically sampling rays, and more. With our renderer, this flexibility comes at little cost compared to rasterization. Experiments demonstrate the speed and accuracy of our approach, as well as several applications in computer graphics and vision. We further propose related improvements to the basic Gaussian representation, including a simple use of generalized kernel functions which significantly reduces particle hit counts.
@@
submissionid: 536copyright: rightsretainedjournal: TOGjournalyear: 2024journalvolume: 43journalnumber: 6article: 232publicationmonth: 12doi: 10.1145/3687934submissionid: 536ccs: Computing methodologies Renderingccs: Computing methodologies Reconstruction
Figure 1. We propose a method for fast forward and inverse ray tracing of particle-based scene representations such as Gaussians. The main idea is to construct encapsulating primitives around each particle, and insert them into a BVH to be rendered by a ray tracer specially adapted to the high density of overlapping particles. Efficient ray tracing opens the door to many advanced techniques, including secondary ray effects like mirrors, refractions and shadows, as well as highly-distorted cameras with rolling shutter effects and even stochastic sampling of rays. Project page: GaussianTracer.github.io
@@ 1. Introduction
Figure 2. Runtime Ray Tracing Effects: Our ray-based pipeline is easily compatible with conventional ray-based visual effects at test time, including reflections (top left), depth of field (top middle), refractions (bottom left), hard shadows cast by meshes (bottom middle), and myriad combinations of them (right).
@@ 2.1. Novel-View Synthesis and Neural Radiance Fields
Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020) have revolutionized the field of novel-view synthesis by representing the scenes in terms of a volumetric radiance field encoded in a coordinate-based neural network. This network can be queried at any location to return the volumetric density and view-dependent color. The photo-realistic quality of NeRFs has made them the standard representation for novel-view synthesis. Follow-up works have focused on improving the speed (Müller et al., 2022; Reiser et al., 2021), quality (Barron et al., 2021, 2022, 2023), and surface representation (Wang et al., 2021; Yariv et al., 2021; Li et al., 2023; Wang et al., 2023a). NeRF has also been extended to large-scale scenes (Turki et al., 2022; Li et al., 2024a), sparse inputs views (Niemeyer et al., 2022), in-the-wild image collections (Martin-Brualla et al., 2021), and reflections (Guo et al., 2022). Finally, several works investigated ways to speed up the inference by baking the neural fields to more performant representations (Reiser et al., 2023; Wang et al., 2023b; Duckworth et al., 2023; Reiser et al., 2024). While achieving high quality and fast rendering speeds, these methods often employ compute-expensive multi-stage training procedures.
@@ 2.2. Point-Based and Particle Rasterization
Jointly, these works have made significant progress, but they still inherit limitations of rasterization. Indeed, they are not able to represent highly distorted cameras, model secondary lighting effects, or simulate sensor properties such as rolling shutter or motion blur. Several works have tried to work around these limitations.  Niemeyer et al. (2024) first train a Zip-NeRF (Barron et al., 2023) that can model distorted and rolling shutter cameras and then render perfect pinhole cameras from the neural field and distill them into a 3DGS representation. This allows for fast inference, but is still limited to perfect pinhole cameras. To address secondary lighting effects, recent works bake occlusion information into spherical harmonics at each Gaussian (Gao et al., 2023; Liang et al., 2023) or leverage shading models and environment maps (Jiang et al., 2024). The latter two of these render only with rasterization; in contrast Gao et al. (2023) traces rays for initial visibility determination, but uses only a visibilty forward pass, and restricts ray tracing to the training phase, relying on rasterization during inference and inheriting its limitations. In contrast, our method uses optimized ray-tracing as the sole renderer throughout both training and inference, which allows for inserting objects, refraction, lens distortion, and other complex effects. Additionally, Gao et al. (2023) use axis-aligned bounding boxes (AABBs) to enclose particles, which results in approximately 3×3$\times$3 × lower FPS during inference compared to the stretched icosahedrons employed in our optimized tracer (Section 5.2). Finally, for complex lens effects, Seiskari et al. (2024) model the motion blur and rolling shutter of the camera by approximating them in screen space through rasterization with pixel velocities. Unlike these works, we formulate a principled method for efficient ray tracing of volumetric particles, which natively alleviates all the limitations mentioned above and further allows us to simulate effects such as depth of field and perfect mirrors.
@@ 2.3. Differentiable Ray Tracing of Volumetric Particles
Our formulation takes root in these precursory works. However as opposed to (Knoll et al., 2019), it is guaranteed to process every particle intersecting the ray, and contrary to (Brüll and Grosch, 2020) the hit processing order is consistent, which ensures the differentiability of our tracing algorithm.
Compared to rasterization, differentiable ray tracing of semi-transparent particles has seen much less progress in recent years. Perhaps the most similar rendering formulation to ours was proposed in Fuzzy Metaballs (Keselman and Hebert, 2022, 2023), but it is limited to scenes with a small set of 3D Gaussian particles (several tens) and images with very low resolution. Different to Fuzzy Metaballs, our method can easily handle scenes with several millions of particles from which it can render full HD images in real-time. In another direction, Belcour et al. (2013) incorporate defocus and motion blur in to ray tracers by leveraging Gaussian approximations as a sampling technique, rather than a scene representation as used here.
@@ 3.2. Differentiable Rendering of Particle Representations
Figure 4. Proxy Geometries: Examples of BVH primitives considered.
@@ 4.1. Bounding Primitives
(7) $\bm{v}\leftarrow\sqrt{2\log(\sigma/\alpha_{\textrm{min}})}\bm{S}\bm{R}^{T}\bm{v}+\bm{\mu}.$
@@ 4.2. Ray Tracing Renderer
Our renderer builds on past approaches for tracing semi-transparent surfaces or particles:  Knoll et al. (2019) repeated gather slabs of particles and sort within each slab, while  Brüll and Grosch (2020) process all semi-transparent surfaces into a $k$-buffer, merging adjacent particles when the list overflows. As discussed in Section 2.3, because of their approximations, these algorithms do not produce a consistent rendering, which prevents differentiation and generates artifacts.
Figure 5. Rendering: on each round of tracing, the next $k$ closest hit particles are collected and sorted into depth order along the ray, the radiance is computed in-order, and the ray is cast again to collect the next hits.
@@
Figure 5, Figure 3, Procedure 4.4, and Procedure 4.4 summarize our approach. To compute incoming radiance along each ray, a ray-gen program traces a ray against the BVH to gather the next $k$ particles, using an any-hit program to maintain a sorted buffer of their indices. For efficiency, at this stage the particle response is not yet evaluated; all primitive hits are treated as intersected particles. The ray-gen program then iterates through the sorted array of primitive hits, retrieves the corresponding particle for each, and renders them according to Equation 6. The process then repeats, tracing a new ray from the last rendered particle to gather the next $k$ particles. The process terminates once all particles intersecting the ray are processed, or early-terminates as soon as enough particle density is intersected to reach a predefined minimum transmittance $T_{\textrm{min}}$. Compared to past approaches this renderer allows for processing the intersection in a consistent order, without missing any particle nor approximating the transmittance.
@@ 4.3. Evaluating Particle Response
4.4. Differentiable Ray Tracing and Optimization
Differentiable Rendering
Beyond forward-rendering of particle scenes, our ray tracing renderer is also differentiable, to support optimizing particle scenes from observed data. To backpropagate (i.e., reverse-mode differentiate) through the renderer with respect to particle parameters, we first perform an ordinary forward-pass render and compute the desired objective functions. Then, in the backward pass we re-cast the same rays to sample the same set of particles in order, computing gradients with respect to each shading expression and accumulating gradients in shared buffers with atomic scatter-add operations. We implemented all derivative expressions by hand in an NVIDIA OptiX ray-gen program which is structurally similar to  Procedure 4.4.
@@ 4.4. Differentiable Ray Tracing and Optimization
5:     if $h.\tau_{\textrm{hit}}<\mathcal{H}[i].\tau_{\textrm{hit}}$ then
@@ 4.5. Particle Kernel Functions
  • Cosine wave modulations along axis $i$:
    (10) $\hat{\rho}_{c}(\bm{x})=\hat{\rho}(\bm{x})(0.5+0.5cos(\psi(\bm{S}^{-1}\bm{R}^{T}(\bm{x}-\bm{\mu}))_{i}))$
@@
MipNeRF360 Tanks & Temples Deep Blending
[…]
Plenoxels 23.63 0.670 - 2.1GB 21.08 0.719 - 2.3GB 23.06 0.795 - 2.7GB
INGP-Base 26.43 0.725 - 13MB 21.72 0.723 - 13MB 23.62 0.797 - 13MB
INGP-Big 26.75 0.751 - 48MB 21.92 0.745 - 48MB 24.96 0.817 - 48MB
MipNeRF360 29.23 0.844 - 8.6MB 22.22 0.759 - 8.6MB 29.40 0.901 - 8.6MB
Zip-NeRF 30.38 0.883 0.197 - - - - - - - - -
3DGS (paper) 28.69 0.871 - 734MB 23.14 0.853 - 411MB 29.41 0.903 - 676MB
3DGS (checkpoint) 28.83 0.867 0.224 763MB 23.35 0.837 - 422MB 29.43 0.898 - 704MB
Ours (reference) 28.69 0.871 0.220 387MB 23.03 0.853 0.193 519MB 29.89 0.908 0.303 329MB
Ours 28.71 0.854 0.248 383MB 23.20 0.830 0.222 489MB 29.23 0.900 0.315 287MB
@@ 5. Experiments and Ablations
In the experiments that follow, we will refer to two variants of our method. The Ours (reference) variant corresponds to (Kerbl et al., 2023) as closely as possible. It employs regular 3D Gaussian particles, and leaves the optimization hyperparameters unchanged. We treat this as a high-quality configuration. The Ours variant is adapted based on the experiments that follow, improving runtime speed at a slight expense of quality. It uses degree-2 generalized Gaussian particles, a density learning rate to $0.09$ during optimizing, as well as optimizing with incoherent rays in a batch size of $2^{19}$ starting after 15,000 training iterations. Empirically, we find that the larger density learning rate of this model produces denser particles. When combined with the faster fall-off of degree-2 generalized Gaussian particles compared to regular Gaussians, this leads to fewer hits along the rays and faster rendering speeds with minimal quality loss.
@@ 5.1.1. MipNeRF360
Table 1 shows quantitative results, while novel-views are qualitatively compared in Figure 7. In terms of quality, our method performs on par or slightly better than 3DGS (Kerbl et al., 2023) and other state-of-the-art methods. For this dataset, we also compare our method against the recent top-tier method of Zip-NeRF (Barron et al., 2023) which achieves 30.38 PSNR. In terms of runtime (Table 2), at 78 FPS our efficient ray tracing implementation is approximately three times slower than rasterization (238 FPS), while maintaining interactive speeds compared to high-quality ray-marching based works MipNeRF360 and Zip-NeRF (¡1 FPS). Zip-NeRF employs multisampling to approximate conical frustums in the ray-casting process, combining MipNeRF360’s anti-aliasing techniques with the speedup mechanism of INGP. While achieving unprecedented rendering quality, Zip-NeRF does not support real-time rendering ($<1$ FPS).
@@ 5.1.4. NeRF Synthetic
5.1.5. Zip-NeRF and Distortion
The rasterization approach in 3DGS (Kerbl et al., 2023) implicitly treats supervision images as coming from perfect-pinhole cameras, whereas our ray-tracing approach can easily be applied directly to highly-distorted cameras such as fisheye captures. Images can be undistorted through postprocessing to enable fitting with 3DGS, but this comes at a cost, including significant cropping or wasted space in the image plane.
We demonstrate this effect on the Zip-NeRF dataset (Barron et al., 2023), which includes four large-scale scenes featuring both indoor and outdoor areas. These scenes are originally captured from highly distorted fisheye cameras, with undistorted versions also provided through postprocessing. Table 3 and Figure 8 compare the quality of training 3DGS on undistorted views, vs. our ray-traced method on the undistorted views or the original distorted images. Our model achieves the highest quality when trained on the original distorted images, partly due to the loss of input supervision signals caused by the cropping of marginal pixels during undistortion. Note that rendering 3DGS from the original fisheye views is impossible, as its tile-based rasterizer is designed for perfect-pinhole rendering.
Figure 8. Qualitative comparison of our results against 3DGS (Kerbl et al., 2023) when trained on distorted or undistorted views and then rendered from undistorted views.
@@ 5.2.1. Particle Primitives
Table 3. Comparison of PSNR achieved by our method versus 3DGS (Kerbl et al., 2023) when trained and tested on distorted or undistorted views.
Test views
Method undistorted original (fisheye)
3DGS w/ undistorted inputs 24.18 N/A
Ours (reference) w/ undistorted inputs 24.59 23.69
Ours (reference) w/ original inputs 24.71 24.40
@@ 5.2.2. Tracing Algorithms
  • Slab tracing (Knoll et al., 2019) (SLAB): tracing slabs along the ray, order independent collection of the $k$-first hits in the any-hit program, sorting and integrating the hits in the ray-gen program.
@@
For each algorithm with a choice of parameters (size of the array, number of slabs, or number of samples), we perform a parameter sweep and present the best-performing setting.
@@ 6.1. Ray-Based Effects
First, we the radiance field rendering pipeline with a variety of visual effects which are naturally compatible with ray tracing (Figure 2 and  13). Here we demonstrate only manually-specified forward rendering, although inverse rendering in concert with these effects is indeed supported by our approach, and is a promising area for ongoing work.
@@
Optical ray effects are supported by interleaved tracing of triangular faces and Gaussian particles. Precisely, we maintain an extra acceleration structure consisting only of mesh faces for additional inserted geometry. When casting each ray in Procedure 4.4, we first cast rays against inserted meshes; if a mesh is hit, we render all particles only up to the hit, and then compute a response based on the material. For refractions and reflections, this means continuing tracing along a new redirected ray according to the laws of optics. For non-transparent diffused meshes, we compute the color and blend it with the current ray radiance, then terminate the ray.
@@
Following (Müller et al., 2022), we simulate depth of field by progressively tracing multiple independent ray samples per pixel (spp), weighted together with a moving average to denoise the output image. The examples in Figures  2 and 13 use 64-256 spp, although convergence is often reached with fewer samples by selecting subsamples with low discrepancy sequences (Burley, 2020).